Volcanic activity at the Soufriere Hills volcano has been at a moderate level today. The activity is still dominated by small to moderate-sized rockfalls from the lava dome. These are concentrated on the northwest side of the dome and confirm the recent change in the focus of activity.
Generally small rockfall signals continue to dominate seismicity at the volcano. There were 54 of these events during this reporting period, identical to yesterday's count. The largest rockfalls were at 11:21 and 13:48 on 10 June. There were two hybrid events today and 15 long period earthquakes. There was one VT earthquake, located east of South Soufriere Hills at a depth of about 6 km. There have been brief periods of intermittent low-amplitude broadband tremor throughout the reporting period. One teleseismic earthquake was recorded by all seismic stations at 00:17 on 10 June.
Clear conditions during most of the night allowed viewing of incandescent material on the northwestern and northeastern flanks of the volcano. Near-continuous streams of hot ash and larger rock fragments were observed from the northwestern flank of the dome and were clearly visible from several areas to the west of the volcano. There were other areas of incandescence visible on the northern and northeastern parts of the dome but these were quite inactive with no rockfalls. Observations of the dome were made under poor viewing conditions from the helicopter this morning. The main point of note was that the moat area behind the Gages Wall was filled almost to the rim with debris and that some larger blocks had already spilled over into upper Fort Ghaut. One of the spines had collapsed and is now lying on the lower slope of the dome between the Farrell's and Gages Walls. There were many small rockfalls on the northern and northwestern flanks of the dome, and very few on the eastern side.
EDM measurements were made today on the eastern and western triangles. There has been about 5 mm of movement of the Castle Peak reflector since yesterday; this is consistent with results over the last week. Lines in the western triangle show movements of less than 1 mm since the last occupation on 16 May.
COSPEC measurements of the SO2 gas concentration in the volcanic plume were carried out again today. The measurements are currently being processed, but the SO2 flux was at a very low level, below 100 tonnes per day.
Richie Robertson left Montserrat this morning after serving as Head of MVO for the last four weeks. LLoyd Lynch arrived on Sunday and is now the Head of MVO.
Scientists at the Montserrat Volcano Observatory are very concerned about the current state of the volcano and the dangers to people and property on both the eastern and upper western flanks. Visits to the evacuated zone should be kept to an absolute minimum. The Tar River, Long Ground and Whites areas remain extremely dangerous. The upper Fort Ghaut and Gages Village areas are now also extremely dangerous and the road between Lee's Estate and Ryners Village has been closed. People should not enter these areas under any circumstances. If they do, they put themselves at direct risk of serious injury or hideous death.